# MedGRPO: Multi-Task Reinforcement Learning for Heterogeneous Medical Video Understanding

### v1

@@ Abstract
Large vision-language models struggle with medical video understanding, where spatial precision, temporal reasoning, and clinical semantics are critical. To address this, we first introduce MedVidBench, a large-scale benchmark of 531,850 video-instruction pairs across 8 medical sources spanning video, segment, and frame-level tasks, curated through a rigorous quality assurance pipeline with expert-guided prompting and dual-model validation. While supervised fine-tuning on MedVidBench yields noticeable gains, standard Reinforcement Learning (RL) fails due to imbalanced reward scales across datasets, which destabilizes optimization and leads to training collapse. To overcome this, we introduce MedGRPO, a novel RL framework for balanced multi-dataset training with two key innovations: (1) cross-dataset reward normalization that maps each dataset’s median performance to a common reward value, ensuring fair optimization regardless of difficulty, and (2) a medical LLM judge that evaluates caption quality on five clinical dimensions through comparative similarity scoring. Supervised fine-tuning Qwen2.5-VL-7B on MedVidBench substantially outperforms GPT-4.1 and Gemini-2.5-Flash across all tasks, demonstrating MedVidBench’s efficacy, while our MedGRPO framework further improves upon the SFT baseline across grounding and captioning tasks. Our work establishes a foundational benchmark and robust training methodology for advancing vision-language models in medical domains. Our project website is available at https://yuhaosu.github.io/MedGRPO/.
@@ 1 Introduction
Large vision-language models (VLMs) have demonstrated remarkable capabilities in understanding and reasoning about general-domain visual content [flamingo, blip2, llava, qwen2vl]. However, their performance significantly degrades when applied to high-stake, expert-driven domains such as medicine. Medical video understanding presents unique challenges: it requires interpreting fine-grained surgical actions, understanding domain-specific terminology (e.g., correctly identifying a “grasper” rather than generic “tool”), assessing procedural safety, and reasoning about multi-phase temporal workflows—requirements far beyond the everyday activities on which most models are trained.
@@
Current state-of-the-art VLMs, including GPT-4.1 [gpt4v], Gemini-2.5-Pro [gemini25], and Qwen2.5-VL [qwen2.5vl], struggle to provide adequate performance on medical video tasks. The primary bottleneck is the lack of suitable training data in instruction-following format. While existing medical video datasets such as CholecT50 [cholect50], EgoSurgery [egosurgery], AVOS [avos], and NurViD [nurvid] contain rich annotations, ranging from frame-wise bounding boxes and action triplets to phase labels and audio, they are not designed for VideoLLM which requires question-answer pairs in conversational format. Recent efforts like SurgLaVi [surglavi] and SurgLLM [surgllm] remain limited in task diversity and primarily focus on surgical procedures. Converting existing medical video annotations into high-quality instruction-following QA pairs at scale presents a fundamental challenge: medical annotations require expert-level understanding to transform into natural language that preserves clinical accuracy.
@@
We introduce MedVidBench, a large-scale instructional dataset comprising 532K video-instruction pairs across 8 medical sources and 8 diverse tasks spanning three temporal granularities: video-level understanding (summarization, critical view of safety, next action prediction, skill assessment), segment-level reasoning (temporal action grounding, dense captioning, region captioning), and frame-level grounding (spatiotemporal localization). The core innovation lies in our multi-perspective quality assurance pipeline that systematically converts existing expert annotations into instruction-following format. Rather than creating annotations from scratch, we leverage GPT-4.1 and Gemini-2.5-Flash to transform existing annotations—frame-wise bounding boxes, action triplets, phase labels—into natural language QA pairs. Our pipeline employs source-specific strategies: for datasets with dense frame-wise annotations (CholecT50, EgoSurgery), we use visual prompting by overlaying bounding boxes and labels on frames; for web-sourced videos (AVOS, NurViD), we extract high-quality audio using Whisper-X [whisperx] and enrich prompts with video metadata (e.g., video titles). Dual-model validation ensures consistency and reduces generation errors.
@@
While Supervised Fine-Tuning (SFT) a VLM on MedVidBench establishes a strong baseline, naively applying standard RL algorithms, such as GRPO [deepseekmath], to our heterogeneous dataset to improve performance reveals a critical failure mode: training collapse. Models rapidly overfit to easy datasets—for instance, CoPESD [copesd] achieves median spatiotemporal grounding mIoU around 0.5—while performance on challenging datasets like EgoSurgery [egosurgery] (median mIoU around 0.12) degrades dramatically. This occurs because raw task metrics used as rewards create fundamentally unfair optimization: the model receives consistently higher rewards for easy dataset samples, causing gradient updates to prioritize easy sources while destabilizing learning on hard ones. A second challenge emerges in caption generation tasks: standard semantic similarity metrics fail to capture medical correctness. Two captions may achieve high surface-level similarity yet differ critically in instrument specificity (“tool” vs. “grasper”), action precision (“grasps” vs. “dissects”), anatomical accuracy (“tissue” vs. “cystic duct”), and spatial detail (“upper area” vs. “upper right quadrant”).
@@
We first train Qwen2.5-VL-7B [qwen2.5vl] using supervised fine-tuning (SFT) on MedVidBench and outperform closed-source alternatives including GPT-4.1 and Gemini-2.5-Flash. We then use reinforcement learning with MedGRPO to train this SFT baseline and achieve substantial improvements over the SFT baseline across nearly all tasks while maintaining balanced learning across both easy and hard datasets. Our main contributions are:
@@ 2 Related Work
Medical Video Datasets. Medical video understanding relies on specialized datasets capturing clinical procedures. Surgical datasets like CholecT50 [cholect50] and EgoSurgery [egosurgery] provide frame-level annotations including action triplets and phase labels, but are limited in scale and diversity. Web-sourced datasets like AVOS [avos] and NurViD [nurvid] offer greater diversity with phase labels from which we extract audio, but lack frame-level precision. However, existing datasets provide annotations for traditional computer vision tasks rather than conversational question-answering format required by VideoLLMs, and remain isolated by procedure type. Our work systematically converts existing expert annotations into instruction-following format and enables joint training across 8 heterogeneous sources through reward normalization.
@@
General Video LLMs. Large vision-language models have demonstrated impressive capabilities in general video understanding. Early frameworks like Flamingo [flamingo] and BLIP-2 [blip2] established cross-modal alignment on images, with recent VideoLLMs like Video-ChatGPT [videochatgpt], VideoLLaMA [videollama3], and Video-LLaVA [videollava] extending these to videos through temporal modeling and video instruction tuning. State-of-the-art models including GPT-4.1 [gpt4v], Gemini-2.5-Flash [gemini25], and Qwen2.5-VL [qwen2.5vl] demonstrate strong performance on general-domain benchmarks [activitynet, msrvtt]. However, medical video understanding presents unique challenges: precise instrument identification, fine-grained action recognition, anatomical structure localization, and temporal reasoning about multi-step procedures.
@@
Medical VideoLLMs. Medical video understanding requires specialized knowledge beyond general vision-language models. Early medical VLMs focused on static radiology images (MedVInT [medvint], LLaVA-Med [llavamed], Med-Flamingo [medflamingo]), while recent work addresses RGB surgical videos with temporal dynamics. SurgLaVi [surglavi] introduces 240K surgical clip-caption pairs for CLIP-style contrastive pretraining but remains in descriptive format rather than instruction-following. Building on this, SurgLLM [surgllm] fine-tunes Qwen2-VL on surgical instruction data from Cholec80 [cholec80], a laparoscopic cholecystectomy dataset, demonstrating improved phase recognition and captioning. However, these approaches rely on single-dataset training, limiting generalization across diverse procedures. Existing medical VideoLLMs face three key challenges: single-dataset training limiting cross-procedure generalization, narrow task coverage (primarily captioning and phase recognition), and lack of mechanisms to address difficulty disparity in multi-dataset learning. Our work introduces a unified framework training across 8 datasets and 8 tasks through cross-dataset reward normalization, enabling balanced learning across heterogeneous medical video sources.
@@
Multi-Task Learning in VLMs. Multi-task learning leverages task synergies to improve generalization in vision-language models. Works like Unified-IO [unifiedio] and OFA [ofa] demonstrate that single transformers can handle diverse modalities through task-specific prompting. Multi-dataset training introduces optimization challenges including catastrophic forgetting [gem, packnet] and requires careful gradient management. Recent work applies RL to align VLMs with preferences [rlhfv, silkie], but focuses on single-domain or single-task settings. Existing approaches either require architectural modifications (separate task heads [unifiedio], task-specific encoders [packnet], or task prompts [ofa]) or assume homogeneous difficulty distributions. In medical video understanding, no prior work addresses the challenge we identify: when datasets have vastly different difficulty (e.g., median performance 0.5 vs 0.12), standard RL causes training collapse due to reward magnitude disparities. Our cross-dataset reward normalization provides a simple yet effective solution without architectural changes, enabling balanced multi-dataset learning.
@@
Evaluation Metrics for Video Captioning. Automatic evaluation of video captioning has evolved from n-gram metrics [bleu, meteor, cider] to embedding-based approaches like BERTScore [bertscore] and CLIPScore [clipscore]. Recent LLM-as-judge methods [geval, gptscore, prometheus, gavie, lee2024vhelm] show correlation with human judgment. However, medical domain evaluation remains underexplored, and standard metrics fail to distinguish critical clinical differences like instrument specificity (‘tool’ vs. ‘grasper’), action precision (‘grasps’ vs. ‘dissects’), and anatomical accuracy (‘tissue’ vs. ‘cystic duct’). We introduce a medical LLM judge using comparative evaluation to assess caption quality on five clinical dimensions, combined with semantic similarity for robust RL training.
@@ 3.1 MedVidBench: Multi-Granular Medical Video Benchmark
| Dataset | Format | Domain | Tasks | Scale | Quality Assurance |
| SurgLaVi [surglavi] | CLIP-style | Surgery | 1 | 240K | Dual-model |
| Surg-396K [endochat] | Image-QA | Surgery | 5 | 396K | Single model |
| SVU-31K [surgvidlm] | Video-QA | Surgery | 3 | 31K | Single model |
| MedVidBench | Video-QA | Surgery | 8 | 532K | Dual-model |
| (Ours) | & Instructional | validation |
@@
Data Sources and Task Coverage. MedVidBench integrates 8 diverse medical video sources covering laparoscopic surgery (CholecT50 [cholect50], CholecTrack20 [cholectrack20], Cholec80-CVS [cholec80cvs], CoPESD [copesd]), open surgery (AVOS [avos], EgoSurgery [egosurgery]), robotic surgery (JIGSAWS [JIGSAWS]), and nursing procedures (NurViD [nurvid]). Our benchmark spans 626 unique videos with duration range 20s–1800s and adaptive FPS 0.1–3.0 optimized per task. We design tasks across three temporal granularities: 1) Video-level tasks include Video Summarization (VS), Critical View of Safety (CVS), Next Action Prediction (NAP), and Skill Assessment (SA), requiring holistic understanding of entire procedures; 2) Segment-level tasks include Temporal Action Grounding (TAG), Dense Video Captioning (DVC), and Region Captioning (RC), enabling event-level reasoning; 3) Frame-level tasks include Spatiotemporal Grounding (STG) for precise instrument localization. This multi-granular design enables models to understand when actions occur (temporal), where instruments are located (spatial), and how procedures should be performed (procedural context).
@@
Multi-Perspective Quality Assurance Pipeline. Transforming existing medical annotations into instruction-following format requires expert-level precision (e.g., distinguishing “Maryland dissector” from generic “tool”). We develop a three-stage pipeline: (1) Expert annotation prompting: leverage dataset-specific strategies—for frame-annotated datasets (e.g., CholecT50, EgoSurgery), we overlay bounding boxes and labels directly on frames with procedure-specific context; for web-sourced datasets (AVOS, NurViD), we extract transcripts from high-quality audio using Whisper-X [whisperx] and enrich with metadata; (2) Dual-model generation: generate captions independently with GPT-4.1 and Gemini-2.5-Flash, using both models for video summary and dense captioning tasks to prevent model-specific biases; (3) Quality validation: compute caption similarities between GPT-4.1 and Gemini-2.5-Flash outputs using sentence-transformers [sentencebert], filter low-quality pairs (similarity 0.3), sample videos with 50–180 frames using adaptive FPS (0.1–3.0) suitable for different task requirements, and create video-level train/test splits (test ratio 0.15). This rigorous pipeline produces 531,850 high-quality instances (461,413 training, 70,437 test).
@@
Human Validation Study. We conduct user studies with medical professionals to empirically validate our expert annotation prompting approach. Medical experts compare captions generated using our annotation-enriched prompts (with overlaid bounding boxes, procedure context, and enriched transcripts) versus baseline generation from raw video frames only, rating on clinical accuracy and terminology precision. Results demonstrate significant preference for annotation-enriched generation, confirming that expert annotation prompting produces superior medical video QA pairs compared to naive frame-based generation. More dataset statistics, data curation instruction prompts, LLM-judge prompts, and human study detail in Supplementary.

### v2

@@ Abstract
Large vision-language models struggle with medical video understanding, where spatial precision, temporal reasoning, and clinical semantics are critical. To address this, we first introduce MedVidBench, a large-scale benchmark of 531,850 video-instruction pairs across 8 medical sources spanning video, segment, and frame-level tasks, curated through a rigorous quality assurance pipeline with expert-guided prompting and dual-model validation. While supervised fine-tuning on MedVidBench yields noticeable gains, standard Reinforcement Learning (RL) fails due to imbalanced reward scales across datasets, which destabilizes optimization and leads to training collapse. To overcome this, we introduce MedGRPO, a novel RL framework for balanced multi-dataset training with two key innovations: (1) cross-dataset reward normalization that maps each dataset’s median performance to a common reward value, ensuring fair optimization regardless of difficulty, and (2) a medical LLM judge that evaluates caption quality on five clinical dimensions through comparative similarity scoring. Supervised fine-tuning Qwen2.5-VL-7B on MedVidBench outperforms GPT-4.1 and Gemini-2.5-Flash across all tasks, while MedGRPO further improves the SFT baseline on grounding and captioning. Our work establishes a foundational benchmark and training methodology for advancing medical video understanding with VLMs. Our project website is available at: https://gaozhongpai.github.io/MedGRPO-Page/.
@@ 1 Introduction
Large vision-language models (VLMs) have demonstrated remarkable capabilities in understanding and reasoning about general-domain visual content [2, 26, 30, 54]. However, their performance significantly degrades when applied to high-stake, expert-driven domains such as medicine. Medical video understanding presents unique challenges beyond general video understanding problems such as temporal modeling [46] and object grounding [47]: it requires interpreting fine-grained surgical actions, understanding domain-specific terminology (e.g., correctly identifying a “grasper” rather than generic “tool”), assessing procedural safety, and reasoning about multi-phase temporal workflows—requirements far beyond the everyday activities on which most models are trained.
@@
Current state-of-the-art VLMs, including GPT-4.1 [19], Gemini-2.5-Pro [11], and Qwen2.5-VL [4], struggle to provide adequate performance on medical video tasks. The primary bottleneck is the lack of suitable training data in instruction-following format. While existing medical video datasets such as CholecT50 [40], EgoSurgery [14], AVOS [16], and NurViD [36] contain rich annotations, ranging from frame-wise bounding boxes, action triplets to phase labels, they are not designed for VideoLLM which requires question-answer pairs in conversational format. Recent efforts like SurgLaVi [42] and SurgLLM [10] remain limited in task diversity and primarily focus on surgical procedures. Converting existing medical video annotations into high-quality instruction-following QA pairs at scale presents a fundamental challenge: medical annotations require expert-level understanding to transform into natural language that preserves clinical accuracy.
@@
We introduce MedVidBench, a large-scale instructional dataset comprising 532K video-instruction pairs across 8 medical sources and 8 diverse tasks spanning three temporal granularities: video-level understanding (summarization, critical view of safety, next action prediction, skill assessment), segment-level reasoning (temporal action grounding, dense captioning, region captioning), and frame-level grounding (spatiotemporal localization). The core innovation lies in our multi-perspective quality assurance pipeline that systematically converts existing expert annotations into instruction-following format. Rather than creating annotations from scratch, we leverage GPT-4.1 and Gemini-2.5-Flash to transform existing annotations—frame-wise bounding boxes, action triplets, phase labels—into natural language QA pairs. Our pipeline employs source-specific strategies: for datasets with dense frame-wise annotations (CholecT50, EgoSurgery), we use visual prompting by overlaying bounding boxes and labels on frames; for web-sourced videos (AVOS, NurViD), we extract high-quality transcripts using Whisper-X [5] and enrich prompts with video metadata (e.g., video titles). Dual-model validation ensures consistency and reduces generation errors.
@@
While Supervised Fine-Tuning (SFT) a VLM on MedVidBench establishes a strong baseline, naively applying standard RL algorithms, such as GRPO [45], to our heterogeneous dataset to improve performance reveals a critical failure mode: training collapse. Models rapidly overfit to easy datasets—for instance, CoPESD [53] achieves median spatiotemporal grounding mIoU around 0.5—while performance on challenging datasets like EgoSurgery [14] (median mIoU around 0.12) degrades dramatically. This occurs because raw task metrics used as rewards create fundamentally unfair optimization: the model receives consistently higher rewards for easy dataset samples, causing gradient updates to prioritize easy sources while destabilizing learning on hard ones. A second challenge emerges in caption generation tasks: standard semantic similarity metrics fail to capture medical correctness. Two captions may achieve high surface-level similarity yet differ critically in instrument specificity (“tool” vs. “grasper”), action precision (“grasps” vs. “dissects”), anatomical accuracy (“tissue” vs. “cystic duct”), and spatial detail (“upper area” vs. “upper right quadrant”).
@@
We first train Qwen2.5-VL-7B [4] using supervised fine-tuning (SFT) on MedVidBench and outperform closed-source alternatives including GPT-4.1 and Gemini-2.5-Flash. We then use reinforcement learning with MedGRPO to train this SFT baseline and achieve substantial improvements over the SFT baseline across nearly all tasks while maintaining balanced learning across both easy and hard datasets. Our main contributions are:
@@ 2 Related Work
Medical Video Datasets. Medical video understanding relies on specialized datasets capturing clinical procedures. Surgical datasets like CholecT50 [40] and EgoSurgery [14] provide frame-level annotations including action triplets and phase labels, but are limited in scale and diversity. Web-sourced datasets like AVOS [16] and NurViD [36] offer greater diversity with phase labels from which we extract audio, but lack frame-level precision. However, existing datasets provide annotations for traditional computer vision tasks rather than conversational question-answering format required by VideoLLMs, and remain isolated by procedure type. Our work systematically converts existing expert annotations into instruction-following format and enables joint training across 8 heterogeneous sources through reward normalization.
@@
General Video LLMs. Large vision-language models have demonstrated impressive capabilities in general video understanding. Early frameworks like Flamingo [2] and BLIP-2 [26] established cross-modal alignment on images, with recent VideoLLMs like Video-ChatGPT [34], VideoLLaMA [60], and Video-LLaVA [28] extending these to videos through temporal modeling and video instruction tuning. State-of-the-art models including GPT-4.1 [19], Gemini-2.5-Flash [11], and Qwen2.5-VL [4] demonstrate strong performance on general-domain benchmarks [7, 56]. However, medical video understanding presents unique challenges: precise instrument identification, fine-grained action recognition, anatomical structure localization, and temporal reasoning about multi-step procedures.
@@
Medical VideoLLMs. Medical video understanding requires specialized knowledge beyond general vision-language models. Early medical VLMs focused on static radiology images (MedVInT [62], LLaVA-Med [25], Med-Flamingo [37]), while recent work addresses RGB surgical videos with temporal dynamics. SurgLaVi [42] introduces 240K surgical clip-caption pairs for CLIP-style contrastive pretraining but remains in descriptive format rather than instruction-following. Building on this, SurgLLM [10] fine-tunes Qwen2-VL on surgical instruction data from Cholec80 [49], a laparoscopic cholecystectomy dataset, demonstrating improved phase recognition and captioning. Similarly, Qwen2.5VL-7B [38] fine-tunes on CholecT50 for surgical workflow analyses. However, these approaches rely on single-dataset training, limiting cross-procedure generalization and task coverage, with no mechanisms to handle difficulty disparity across heterogeneous medical video datasets. Concurrent efforts advance medical AI through multimodal reasoning, surgical foundation models, and knowledge distillation [15, 9, 8, 23, 22, 21, 48, 17], yet multi-task learning in medical video understanding remains unexplored. Our work addresses this with a unified framework across 8 datasets and 8 tasks, enabling balanced learning via cross-dataset reward normalization.
@@
Multi-Task Learning in VLMs. Multi-task learning leverages task synergies to improve generalization in vision-language models. Works like Unified-IO [33] and OFA [55] demonstrate that single transformers can handle diverse modalities through task-specific prompting. Multi-dataset training introduces optimization challenges including catastrophic forgetting [32, 35] and requires careful gradient management. Recent work applies RL to align VLMs with preferences [59, 27], but focuses on single-domain or single-task settings. Existing approaches either require architectural modifications (separate task heads [33], task-specific encoders [35], or task prompts [55]) or assume homogeneous difficulty distributions. In medical video understanding, no prior work addresses the challenge we identify: when datasets have vastly different difficulty (e.g., median performance 0.5 vs 0.12), standard RL causes training collapse due to reward magnitude disparities. Our cross-dataset reward normalization provides a simple yet effective solution without architectural changes, enabling balanced multi-dataset learning.
@@
Evaluation Metrics for Video Captioning. Automatic evaluation of video captioning has evolved from n-gram metrics [41, 6, 50] to embedding-based approaches like BERTScore [61] and CLIPScore [18]. Recent LLM-as-judge methods [31, 13, 20, 29, 24] show strong correlation with human judgment, including in clinical settings [12]. However, medical video evaluation remains underexplored, and standard metrics fail to distinguish critical clinical differences like instrument specificity (‘tool’ vs. ‘grasper’), action precision (‘grasps’ vs. ‘dissects’), and anatomical accuracy (‘tissue’ vs. ‘cystic duct’). We introduce a medical LLM judge using comparative evaluation to assess caption quality on five clinical dimensions, combined with semantic similarity for robust RL training.
@@ 3.1 MedVidBench: Multi-Granular Medical Video Benchmark
Data Sources and Task Coverage. MedVidBench integrates 8 diverse medical video sources covering laparoscopic surgery (CholecT50 [40], CholecTrack20 [39], Cholec80-CVS [44], CoPESD [53]), open surgery (AVOS [16], EgoSurgery [14]), robotic surgery (JIGSAWS [1]), and nursing procedures (NurViD [36]). Our benchmark spans 626 unique videos with duration range 20s–1800s and adaptive FPS 0.1–3.0 optimized per task. We design tasks across three temporal granularities: 1) Video-level tasks include Video Summarization (VS), Critical View of Safety (CVS), Next Action Prediction (NAP), and Skill Assessment (SA), requiring holistic understanding of entire procedures; 2) Segment-level tasks include Temporal Action Grounding (TAG), Dense Video Captioning (DVC), and Region Captioning (RC), enabling event-level reasoning; 3) Frame-level tasks include Spatiotemporal Grounding (STG) for precise instrument localization. This multi-granular design enables models to understand when actions occur (temporal), where instruments are located (spatial), and how procedures should be performed (procedural context).
@@
Multi-Perspective Quality Assurance Pipeline. Transforming existing medical annotations into instruction-following format requires expert-level precision (e.g., distinguishing “Maryland dissector” from generic “tool”). We develop a three-stage pipeline: (1) Expert annotation prompting: leverage dataset-specific strategies—for frame-annotated datasets (e.g., CholecT50, EgoSurgery), we overlay bounding boxes and labels directly on frames with procedure-specific context; for web-sourced datasets (AVOS, NurViD), we extract transcripts from high-quality audio using Whisper-X [5] and enrich with metadata; (2) Dual-model generation: generate captions independently with GPT-4.1 and Gemini-2.5-Flash, using both models for video summary and dense captioning tasks to prevent model-specific biases; (3) Quality validation: compute caption similarities between GPT-4.1 and Gemini-2.5-Flash outputs using sentence-transformers [43], filter low-quality pairs (similarity 0.3), sample videos with 50–180 frames using adaptive FPS (0.1–3.0) suitable for different task requirements, and create video-level train/test splits (test ratio 0.15). This rigorous pipeline produces 531,850 high-quality instances (461,413 training, 70,437 test).
@@
Human Validation Study. We conduct user studies with medical professionals to empirically validate our expert annotation prompting approach. Medical experts compare captions generated using our annotation-enriched prompts (with overlaid bounding boxes, procedure context, and enriched transcripts) versus baseline generation from raw video frames only, rating on clinical accuracy and terminology precision. Results demonstrate significant preference for annotation-enriched generation, confirming that expert annotation prompting produces superior medical video QA pairs compared to naive frame-based generation. More dataset statistics, data curation details, LLM-judge prompts, and human study detail in the supplementary.